# Note about non-compact embeddings between Besov spaces

### v1

@@ Abstract.
In this note, we look at the behavior of embedding between Besov spaces and compare its behavior with Sobolev embeddings, mainly when the embeddings are non-compact. We classify that in the case of the non-compact embedding then, depending on parameters, whether or not the embedding is finitely strictly singular and strictly singular. These results could be considered as a continuation and extension of recent results on non-compact “optimal” Sobolev embeddings.
@@ 1. Introduction
(1.1) $W^{k,p}(\Omega)\hookrightarrow L^{p^{*}}(\Omega)\quad\text{and}\quad W^{k,p}(\Omega)\hookrightarrow L^{p^{*},p}(\Omega),$
where $1\leq k<n$, $p\in[1,n/k)$, and $p^{*}=\frac{np}{n-kp}$, are non-compact. Given that $L^{p^{*}}(\Omega)\subsetneq L^{p^{*},p}(\Omega)$, one might expect a significant difference in the “quality” of non-compactness for these embeddings. However, standard measures of non-compactness, such as the decay of entropy numbers or Kolmogorov numbers, fail to differentiate between these embeddings, as both are maximally non-compact (i.e., the measure of non-compactness coincides with the norm of the embeddings, see [LMOP21]).
@@
We provide a comprehensive and complete classification of the “quality” of non-compactness for the embeddings between Besov spaces in the context of strict singularity and finitely strict singularity, as demonstrated in the following two theorems which are one of the main results of this paper.
@@
By applying the standard partition of unity argument, the same results in Theorems 1.1 hold if we replace $\Omega$ by a compact smooth manifold. We leave the details to readers.
@@
We also include a result of embedding of homogeneous Besov spaces ˙$\dot{\mathscr{B}}_{p_{0}q_{0}}^{s_{0}}(\mathbb{R}^{n})\hookrightarrow\dot{\mathscr{B}}_{p_{1}q_{1}}^{s_{1}}(\mathbb{R}^{n})$, see Theorem 5.1.
@@
It is also important to highlight that the Sobolev embedding into the space $L^{p^{*},p}(\Omega)$ is not only the optimal (i.e., smallest) target space within the scale of Lorentz spaces but also the optimal target space among all rearrangement-invariant spaces (such as Lebesgue spaces, Lorentz spaces, Orlicz spaces, and similar spaces). Specifically, this implies that if $Z(\Omega)$ is a rearrangement-invariant space such that $V_{0}^{k,p}(\Omega)\subset Z(\Omega)$, then $L^{p^{*},p}(\Omega)\subset Z(\Omega)$ (see [KP06, Theorem A]). In contrast, the Sobolev embedding into $L^{p^{*}}(\Omega)$ is optimal within the narrower scale of Lebesgue spaces, and the “quality” of non-compactness differs significantly (i.e., non-strictly singular versus finitely strictly singular). This observation suggests that non-strictly singular embeddings may be associated with the “most optimal” embeddings. This insight could help identify cases where we might expect to find the “most optimal” embeddings between Besov spaces, as seen in the case (v) in Theorem 1.1 and case (iii) Theorem 1.2, comparable with cases in which there is a chance for more optimal target spaces, i.e. when embedding is finitely strictly singular or strictly singular.
It is also worth noting that the techniques used in [LM23], [LM19] and [BG89] for Sobolev embeddings are fundamentally different from those employed in this paper. In particular, the results for Sobolev embeddings into $L^{p^{*}}$ require quite sophisticated methods. This contrasts with the techniques used for Besov embeddings and highlights the advantages of working with Besov spaces compared to Sobolev spaces. Additionally, it is interesting to note that, to the best of our knowledge, there is no natural Sobolev embedding that is non-compact, strictly singular, but not finitely strictly singular. This specific case can be observed within the framework of Besov spaces, as demonstrated in Theorem 1.1 and Theorem 1.2, which underscores the richness of Besov spaces.
The paper is structured as follows. In the next section, we recall some definitions and notations used throughout the paper together with the results about wavelet decomposition of Besov spaces and their consequences. Section 3 presents the study of embeddings between sequence spaces. Section 4 contains the proofs of the main theorems, Theorems 1.1 and 1.2. In Section 5 we state and prove the analog of Theorem 1.2 in context of homogeneous Besov spaces, the Theorem 5.1. The paper concludes with Section 6, which contains some final remarks.
@@ 2. Preliminaries
The $n$-th Bernstein number $b_{n}(T)$ of a bounded linear map $T:X\to Y$ is defined by
@@
$T\text{ is compact}\implies T\text{ is finitely strictly singular}\implies T\text{ is strictly singular.}$.
@@
[…]
@@
[…]
@@
Notice that $\ell^{q}(X_{j})_{j=0}^{\infty}$ may not be quasi-normed since if $(X_{j})_{j=0}^{\infty}$ do not have uniform triangle inequality. That is, one may not find a $C>0$ such that $\|x_{j}^{\prime}+x_{j}^{\prime\prime}\|_{X_{j}}\leq C(\|x_{j}^{\prime}\|_{X_{j}}+\|x^{\prime\prime}_{j}\|_{X_{j}})$ for all $j\geq 0$ and $x_{j}^{\prime},x_{j}^{\prime\prime}\in X_{j}$. For example, we can take $X_{j}=\ell^{1/j}$ for $j\geq 0$. However, in our case $(X_{j})_{j=0}^{\infty}$ are always subspaces of $\ell^{p}$ for some $p>0$, where we have $\|x_{j}^{\prime}+x_{j}^{\prime\prime}\|_{X_{j}}\leq 2^{\max(1/p-1,0)}(\|x_{j}^{\prime}\|_{X_{j}}+\|x^{\prime\prime}_{j}\|_{X_{j}})$ for all $j$ and $x_{j}^{\prime},x_{j}^{\prime\prime}$.
@@
For every finite set $S$ and $0<r\leq\infty$, the space (#$(\#S)^{-1/r}\cdot\ell^{r}(S)$ equals to $L^{r}(S,\mu_{S})$ where $\mu_{S}\{a\}=(\#S)^{-1}$ for every $a\in S$.
@@
[…]
@@
We define the closed subspace ~$\widetilde{\mathscr{B}}_{pq}^{s}(\overline{\Omega})\subseteq\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})$ by
~$\widetilde{\mathscr{B}}_{pq}^{s}(\overline{\Omega}):=\{\tilde{f}\in\mathscr{B}_{pq}^{s}(\mathbb{R}^{n}):\operatorname{supp}f\subseteq\overline{\Omega}\}.$
@@
and when $\Omega=\mathbb{R}^{n}$ we have by definition the coincidence of notations
@@
Let $\Omega_{0},\Omega_{1}\subseteq\mathbb{R}^{n}$ be two open subsets such that $\Omega_{0}$ is precompact in $\Omega_{1}$. We have the natural topological embedding ~$\widetilde{\mathscr{B}}_{pq}^{s}(\overline{\Omega}_{0})\hookrightarrow\mathscr{B}_{pq}^{s}(\Omega_{1})$.
@@
On the other hand, take a $\chi\in C_{c}^{\infty}(\Omega_{1})$ such that $\chi|_{\Omega_{0}}\equiv 1$. Recall from e.g. [Tri10, Theorem 2.8.2] that $[\tilde{f}\mapsto\chi\tilde{f}]:\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})\to\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})$ is a bounded map whose kernel contains $\{\tilde{f}:\tilde{f}|_{\Omega_{1}}=0\}$. By the quotient characterization (2.1) it descends to a bounded linear map $[f\mapsto\chi f]:\mathscr{B}_{pq}^{s}(\Omega_{1})\to\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})$, i.e. there is a $C>0$ such that $\|\chi f\|_{\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})}\leq C\|f\|_{\mathscr{B}_{pq}^{s}(\Omega_{1})}$ for all $f\in\mathscr{B}_{pq}^{s}(\Omega_{1})$.
@@
    For every $p,q\in[\varepsilon,\infty],s\in[-\varepsilon^{-1},\varepsilon^{-1}]$ and $f\in\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})$, we have a decomposition (which is orthonormal in $L^{2}$)
@@
    (2.4) $\begin{array}[]{rcl}\Lambda=\Lambda_{\varepsilon}:\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})&\xrightarrow{\simeq}&\ell^{q}\big{(}2^{j(s-\frac{n}{p}+\frac{n}{2})}\cdot\ell^{p}(Q_{j}\times G_{j})\big{)}_{j=0}^{\infty};\\ f&\longmapsto&\{(f,\phi_{jm}^{g})_{L^{2}}:m\in Q_{j},\ g\in G_{j}\}_{j=0}^{\infty};\\ \sum_{j=0}^{\infty}\sum_{m,g}a_{jm}^{g}\psi_{jm}^{g}&\mathrel{\reflectbox{$\longmapsto$}}&\{a_{jm}^{g}:m\in Q_{j},\ g\in G_{j}\}_{j=0}^{\infty}.\end{array}$
@@
Let $\Omega\subset\mathbb{R}^{n}$ be a Lipschitz domain and let $U\supset\overline{\Omega}$ be a bounded open neighborhood. For every $\varepsilon>0$, there is an extension operator $E$ such that $E:\mathscr{B}_{pq}^{s}(\Omega)\to\widetilde{\mathscr{B}}_{pq}^{s}(\overline{U})$ is defined and bounded for all $p,q\in[\varepsilon,\infty]$ and $s\in[-\varepsilon^{-1},\varepsilon^{-1}]$.
@@
Here, we also remark that $E:\mathscr{B}_{pq}^{s}(\Omega)\to\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})$ is always a topological embedding since it is the right inverse to the restriction mapping $[\tilde{f}\mapsto\tilde{f}|_{\Omega}]:\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})\twoheadrightarrow\mathscr{B}_{pq}^{s}(\Omega)$.
@@
    #$\#R_{j}=2^{j(n-A)}$ for $j\geq A$ and we have $\Lambda^{-1}:\ell^{q}(2^{j(s-\frac{n}{p}+\frac{n}{2})}\cdot\ell^{p}(R_{j}))_{j=A}^{\infty}\to\widetilde{\mathscr{B}}_{pq}^{s}(\overline{V})$. As a result $\Lambda^{-1}:\ell^{q}(2^{j(s-\frac{n}{p}+\frac{n}{2})}\cdot\ell^{p}(R_{j}))_{j=A}^{\infty}\to\mathscr{B}_{pq}^{s}(\Omega)$ is a topological embedding.
@@
    #$\#S_{j}=2^{(j+A)n}$ for $j\geq 0$ and we have $\Lambda:\widetilde{\mathscr{B}}_{pq}^{s}(\overline{U})\to\ell^{q}(2^{j(s-\frac{n}{p}+\frac{n}{2})}\cdot\ell^{p}(S_{j}))_{j=0}^{\infty}$. As a result $\Lambda\circ E:\mathscr{B}_{pq}^{s}(\Omega)\to\ell^{q}(2^{j(s-\frac{n}{p}+\frac{n}{2})}\cdot\ell^{p}(S_{j}))_{j=0}^{\infty}$ is a topological embedding.
@@
Recall $Q_{j}=2^{-j}\mathbb{Z}^{n}$, thus for a large enough $A\geq 1$ (depending on $U,V,L$) we have #$\#\mathfrak{R}_{j}\geq 2^{(j-A)n}$ (for $j\geq A$) and #$\#\mathfrak{S}_{j}\leq 2^{(j+A-1)n}$ (for $j\geq 0$).
By Proposition 2.12 ii we have $\operatorname{supp}\psi_{jm}^{g}\subset V$ for all $m\in\mathfrak{R}_{j}$ and $g\in G_{j}$, thus by (2.3) $\sum_{j=A}^{\infty}\sum_{m\in\mathfrak{R}_{j},g\in G_{j}}a_{jm}^{g}\psi_{jm}^{g}\in\widetilde{\mathscr{B}}_{pq}^{s}(\overline{V})$ holds for all $\{a_{jm}^{g}\}_{j=A}^{\infty}\in\ell^{q}(2^{j(s-\frac{n}{p}+\frac{n}{2})}\cdot\ell^{p}(\mathfrak{R}_{j}\times G_{j}))_{j=0}^{\infty}$. Taking $R_{j}\subseteq\mathfrak{R}_{j}\times G_{j}$ be any set such that #$\#R_{j}=2^{(j-A)n}$ and applying Proposition 2.12 iv we get $\Lambda^{-1}:\ell^{q}(2^{j(s-\frac{n}{p}+\frac{n}{2})}\cdot\ell^{p}(R_{j}))_{j=A}^{\infty}\to\widetilde{\mathscr{B}}_{pq}^{s}(\overline{V})$. Applying Lemma 2.11 we have topological embedding ~$\widetilde{\mathscr{B}}_{pq}^{s}(\overline{V})\hookrightarrow\mathscr{B}_{pq}^{s}(\Omega)$. This completes the proof of i.
Similarly by Proposition 2.12 ii and iii, for every $f\in\widetilde{\mathscr{B}}_{pq}^{s}(\overline{U})$, we have $(f,\phi_{jm}^{g})_{L^{2}}=0$ if $m\notin\mathfrak{S}_{j}$, hence by Proposition 2.12 iii $f=\sum_{j=0}^{\infty}\sum_{m\in\mathfrak{S}_{j},g\in G_{j}}(f,\psi_{jm}^{g})_{L^{2}}\cdot\psi_{jm}^{g}$. Taking $S_{j}\supseteq\mathfrak{S}_{j}\times G_{j}$ be any set in $Q_{j}\times G_{j}$ such that #$\#S_{j}=2^{(j+A)n}$ and applying Proposition 2.12 iv we get $\Lambda:\widetilde{\mathscr{B}}_{pq}^{s}(\overline{U})\to\ell^{q}(2^{j(s-\frac{n}{p}+\frac{n}{2})}\cdot\ell^{p}(S_{j}))_{j=0}^{\infty}$. Applying Lemma 2.14 which gives $E:\mathscr{B}_{pq}^{s}(\Omega)\to\mathscr{B}_{pq}^{s}(\overline{U})$, we get the required boundedness of $\Lambda\circ E$. Since $\Lambda$ and $E$ are both topological embeddings, so does $\Lambda\circ E$. This completes the proof of ii. ∎
@@
Indeed one can check that $\ell^{q_{0}}(2^{j(s-\frac{n}{p_{0}}+\frac{n}{2})}\cdot\ell^{p_{0}})_{j=0}^{\infty}\subset\ell^{q_{1}}(2^{j(s-\frac{n}{p_{1}}+\frac{n}{2})}\cdot\ell^{p_{1}})_{j=0}^{\infty}$ holds if and only if either (2.5) holds; and the map is always non-compact. By Proposition 2.12 iv we get i.
Similarly $\ell^{q_{0}}(2^{j(s-\frac{n}{p_{0}}+\frac{n}{2})}\cdot\ell^{p_{0}}\{1,\dots,2^{nj}\})_{j=0}^{\infty}\subset\ell^{q_{1}}(2^{j(s-\frac{n}{p_{1}}+\frac{n}{2})}\cdot\ell^{p_{1}}\{1,\dots,2^{nj}\})_{j=0}^{\infty}$ if and only if either (2.6) or (2.7) holds; and the map is compact if and only if (2.7) holds.

### v2

# A Classification Theorem on Non-compact Embeddings between Besov Spaces

@@ Abstract.
We analyze the embedding properties between Besov spaces, defined on the total space $\mathbb{R}^{n}$ and on bounded domains. We give a complete classification on whether or not these embedding maps satisfy certain weak compactness characterized by the so-called strictly and finitely strictly singular condition. The result extends the recent findings on Sobolev embeddings by offering a refined description of the quality of non‐compactness in the setting of Besov spaces.
@@ 1. Introduction
$W^{k,p}(\Omega)\hookrightarrow L^{\frac{np}{n-kp}}(\Omega)\quad\text{and}\quad W^{k,p}(\Omega)\hookrightarrow L^{\frac{np}{n-kp},p}(\Omega),$
where $1\leq k<n$ and $p\in[1,n/k)$ are non-compact. Given that $L^{\frac{np}{n-kp}}(\Omega)\subsetneq L^{\frac{np}{n-kp},p}(\Omega)$, one might expect a significant difference in the “quality” of non-compactness for these embeddings. However, standard measures of non-compactness, such as the decay of entropy numbers or Kolmogorov numbers, fail to differentiate between these embeddings, as both are maximally non-compact (i.e., the measure of non-compactness coincides with the norm of the embeddings, see [LMOP21]).
@@
We provide a comprehensive and complete classification of the “quality” of non-compactness for the embeddings between Besov spaces in the context of strict singularity and finitely strict singularity, as demonstrated in the following two theorems which are the main results of this paper.
@@
By applying the standard partition of unity argument, the same results in Theorems 1.1 hold if we replace $\Omega$ by a compact smooth manifold. We refer to [Tri92, Chapter 7] for the definition of spaces and leave the details to readers.
@@
We also include a result of embedding of homogeneous Besov spaces $\dot{\mathscr{B}}_{p_{0}q_{0}}^{s_{0}}(\mathbb{R}^{n})\hookrightarrow\dot{\mathscr{B}}_{p_{1}q_{1}}^{s_{1}}(\mathbb{R}^{n})$, see Theorem 5.1.
@@
It is also important to highlight that the Sobolev embedding $W^{k,p}(\Omega)$ into the space $L^{\frac{np}{n-kp},p}(\Omega)$ is not only the optimal (i.e., smallest) target space within the scale of Lorentz spaces but also the optimal target space among all rearrangement-invariant spaces (such as Lebesgue spaces, Lorentz spaces, Orlicz spaces, and similar spaces). Specifically, this implies that if $Z(\Omega)$ is a rearrangement-invariant space such that $V_{0}^{k,p}(\Omega)\subset Z(\Omega)$, then $L^{\frac{np}{n-kp},p}(\Omega)\subseteq Z(\Omega)$ (see [KP06, Theorem A]). In contrast, the Sobolev embedding into $L^{\frac{np}{n-kp}}(\Omega)$ is optimal within the narrower scale of Lebesgue spaces, and the “quality” of non-compactness differs significantly (i.e., non-strictly singular versus finitely strictly singular). This observation suggests that non-strictly singular embeddings may be associated with the “most optimal” embeddings. This insight could help identify cases where we might expect to find the “most optimal” embeddings between Besov spaces, as seen in the case v in Theorem 1.1 and case iii in Theorem 1.2, comparable with cases in which there is a chance for more optimal target spaces, i.e. when embedding is finitely strictly singular or strictly singular.
It is also worth noting that the techniques used in [LM23], [LM19] and [BG89] for Sobolev embeddings are fundamentally different from those employed in this paper. In particular, the results for Sobolev embeddings $W^{k,p}\to L^{\frac{np}{n-kp}}$ require quite sophisticated methods. This contrasts with the techniques used for Besov embeddings and highlights the advantages of working with Besov spaces compared to Sobolev spaces and Lorentz spaces. This is demonstrated in our forthcoming paper [CLY25], in which, by using the results obtained for Besov spaces and their relation with Sobolev spaces, we improved and extend findings from [BG89] and [LM23].
Additionally, it is noteworthy that, to the best of our knowledge, there is no natural Sobolev embedding that is non-compact, strictly singular, yet not finitely strictly singular. This specific scenario, however, can be naturally observed within the framework of Besov spaces, as demonstrated in Theorem 1.1 and Theorem 1.2, underscoring the structural richness of Besov spaces.
The paper is structured as follows. In the next section, we recall essential definitions and notations used throughout the paper, along with results concerning the wavelet decomposition of Besov spaces and their implications. Section 3 focuses on the study of embeddings between sequence spaces. In Section 4, we present the proofs of the main theorems, Theorems 1.1 and 1.2. Section 5 extends Theorem 1.2 to the setting of homogeneous Besov spaces, resulting in Theorem 5.1. Finally, Section 6 concludes the paper with final remarks and observations.
@@ 2. Preliminaries
For $n\geq 1$, the $n$-th Bernstein number $b_{n}(T)$ of a bounded linear map $T:X\to Y$ is
@@
[…]
@@
We also write $x=\sum_{j=0}^{\infty}x_{j}\otimes e_{j}$ where $e_{j}=(0,\dots,0,\underset{j}{1},0,\dots)$ is the standard basis in $\ell^{q}$.
@@
Using the notation in the definition, for the double sequence space $\ell^{q}(\ell^{p})$, by the element $e_{j}\otimes e_{k}$ we mean $e_{j}\in\ell^{p}$ and $e_{k}\in\ell^{q}$. More precisely $e_{j}\otimes e_{k}=(0,\dots,0,\underset{k}{e_{j}},0,\dots)$.
@@
Notice that $\ell^{q}(X_{j})_{j=0}^{\infty}$ may not be quasi-normed since $(X_{j})_{j=0}^{\infty}$ may not have uniform triangle inequality. That is, one may not find a $C>0$ such that $\|x_{j}^{\prime}+x_{j}^{\prime\prime}\|_{X_{j}}\leq C(\|x_{j}^{\prime}\|_{X_{j}}+\|x^{\prime\prime}_{j}\|_{X_{j}})$ for all $j\geq 0$ and $x_{j}^{\prime},x_{j}^{\prime\prime}\in X_{j}$. For example, we can take $X_{j}=\ell^{1/j}$ for $j\geq 0$. However, in our case $(X_{j})_{j=0}^{\infty}$ are always subspaces of $\ell^{p}$ for some $p>0$, where we have $\|x_{j}^{\prime}+x_{j}^{\prime\prime}\|_{X_{j}}\leq 2^{\max(1/p-1,0)}(\|x_{j}^{\prime}\|_{X_{j}}+\|x^{\prime\prime}_{j}\|_{X_{j}})$ for all $j$ and $x_{j}^{\prime},x_{j}^{\prime\prime}$.
@@
For every finite set $S$ and $0<r\leq\infty$, the space $(\#S)^{-1/r}\cdot\ell^{r}(S)$ equals to $L^{r}(S,\mu_{S})$ where $\mu_{S}\{a\}=(\#S)^{-1}$ for every $a\in S$.
@@
Here $(\varphi_{j})_{j=0}^{\infty}\subset\mathscr{S}(\mathbb{R}^{n})$ is a fixed family of Schwartz functions whose Fourier transform $\hat{\varphi}_{j}(\xi)=\int\varphi_{j}(x)e^{-2\pi ix\xi}dx$ satisfy
@@
We define the closed subspace $\widetilde{\mathscr{B}}_{pq}^{s}(\overline{\Omega})\subseteq\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})$ by
$\widetilde{\mathscr{B}}_{pq}^{s}(\overline{\Omega}):=\{\tilde{f}\in\mathscr{B}_{pq}^{s}(\mathbb{R}^{n}):\operatorname{supp}f\subseteq\overline{\Omega}\}.$
Remark 2.11.
@@
When $\Omega=\mathbb{R}^{n}$ we have by definition the coincidence of notations
@@
Let $\Omega_{0},\Omega_{1}\subseteq\mathbb{R}^{n}$ be two open subsets such that $\Omega_{0}$ is precompact in $\Omega_{1}$. We have the natural topological embedding $\widetilde{\mathscr{B}}_{pq}^{s}(\overline{\Omega}_{0})\hookrightarrow\mathscr{B}_{pq}^{s}(\Omega_{1})$.
@@
On the other hand, take a $\chi\in C_{c}^{\infty}(\Omega_{1})$ such that $\chi|_{\Omega_{0}}\equiv 1$. Recall from e.g. [Tri10, Theorem 2.8.2] that $\tilde{f}\mapsto\chi\tilde{f}]:\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})\to\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})$ is a bounded map whose kernel contains $\{\tilde{f}:\tilde{f}|_{\Omega_{1}}=0\}$. By the quotient characterization (2.1) it descends to a bounded linear map $f\mapsto\chi f]:\mathscr{B}_{pq}^{s}(\Omega_{1})\to\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})$, i.e. there is a $C>0$ such that $\|\chi f\|_{\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})}\leq C\|f\|_{\mathscr{B}_{pq}^{s}(\Omega_{1})}$ for all $f\in\mathscr{B}_{pq}^{s}(\Omega_{1})$.
@@
    For every $p,q\in[\varepsilon,\infty],s\in[-\varepsilon^{-1},\varepsilon^{-1}]$ and $f\in\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})$, we have expansion (which is orthonormal in $L^{2}$)
@@
    (2.4) $\begin{array}[]{rcl}\Lambda=\Lambda_{\varepsilon}:\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})&\xrightarrow{\simeq}&\ell^{q}\big(2^{j(s-\frac{n}{p}+\frac{n}{2})}\cdot\ell^{p}(Q_{j}\times G_{j})\big)_{j=0}^{\infty};\\ f&\longmapsto&\{(f,\psi_{jm}^{g})_{L^{2}}:m\in Q_{j},\ g\in G_{j}\}_{j=0}^{\infty};\\ \sum_{j=0}^{\infty}\sum_{m,g}a_{jm}^{g}\psi_{jm}^{g}&\mathrel{\reflectbox{$\longmapsto$}}&\{a_{jm}^{g}:m\in Q_{j},\ g\in G_{j}\}_{j=0}^{\infty}.\end{array}$
@@
Let $\Omega\subset\mathbb{R}^{n}$ be a bounded Lipschitz domain and let $U\supset\overline{\Omega}$ be an open neighborhood. For every $\varepsilon>0$, there is an extension operator $E$ such that $E:\mathscr{B}_{pq}^{s}(\Omega)\to\widetilde{\mathscr{B}}_{pq}^{s}(\overline{U})$ is defined and bounded for all $p,q\in[\varepsilon,\infty]$ and $s\in[-\varepsilon^{-1},\varepsilon^{-1}]$.
@@
Here, we also remark that $E:\mathscr{B}_{pq}^{s}(\Omega)\to\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})$ is always a topological embedding since it is the right inverse to the restriction mapping $\tilde{f}\mapsto\tilde{f}|_{\Omega}]:\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})\twoheadrightarrow\mathscr{B}_{pq}^{s}(\Omega)$.
@@
    $R_{j}=2^{(j-A)n}$ for $j\geq A$ and we have $\Lambda^{-1}:\ell^{q}(2^{j(s-\frac{n}{p}+\frac{n}{2})}\cdot\ell^{p}(R_{j}))_{j=A}^{\infty}\to\widetilde{\mathscr{B}}_{pq}^{s}(\overline{V})$. As a result $\Lambda^{-1}:\ell^{q}(2^{j(s-\frac{n}{p}+\frac{n}{2})}\cdot\ell^{p}(R_{j}))_{j=A}^{\infty}\to\mathscr{B}_{pq}^{s}(\Omega)$ is a topological embedding.
@@
    $S_{j}=2^{(j+A)n}$ for $j\geq 0$ and we have $\Lambda:\widetilde{\mathscr{B}}_{pq}^{s}(\overline{U})\to\ell^{q}(2^{j(s-\frac{n}{p}+\frac{n}{2})}\cdot\ell^{p}(S_{j}))_{j=0}^{\infty}$. As a result $\Lambda\circ E:\mathscr{B}_{pq}^{s}(\Omega)\to\ell^{q}(2^{j(s-\frac{n}{p}+\frac{n}{2})}\cdot\ell^{p}(S_{j}))_{j=0}^{\infty}$ is a topological embedding.
@@
Recall $Q_{j}=2^{-j}\mathbb{Z}^{n}$, thus for a large enough $A\geq 1$ (depending on $U,V,L$) we have $\mathfrak{R}_{j}\geq 2^{(j-A)n}$ (for $j\geq A$) and $\mathfrak{S}_{j}\leq 2^{(j+A-1)n}$ (for $j\geq 0$).
By Proposition 2.13 ii we have $\operatorname{supp}\psi_{jm}^{g}\subset V$ for all $m\in\mathfrak{R}_{j}$ and $g\in G_{j}$, thus by (2.3) $\sum_{j=A}^{\infty}\sum_{m\in\mathfrak{R}_{j},g\in G_{j}}a_{jm}^{g}\psi_{jm}^{g}\in\widetilde{\mathscr{B}}_{pq}^{s}(\overline{V})$ holds for all $\{a_{jm}^{g}\}_{j=A}^{\infty}\in\ell^{q}(2^{j(s-\frac{n}{p}+\frac{n}{2})}\cdot\ell^{p}(\mathfrak{R}_{j}\times G_{j}))_{j=0}^{\infty}$. Taking $R_{j}\subseteq\mathfrak{R}_{j}\times G_{j}$ be any set such that $R_{j}=2^{(j-A)n}$ and applying Proposition 2.13 iv we get $\Lambda^{-1}:\ell^{q}(2^{j(s-\frac{n}{p}+\frac{n}{2})}\cdot\ell^{p}(R_{j}))_{j=A}^{\infty}\to\widetilde{\mathscr{B}}_{pq}^{s}(\overline{V})$. Applying Lemma 2.12 we have topological embedding $\widetilde{\mathscr{B}}_{pq}^{s}(\overline{V})\hookrightarrow\mathscr{B}_{pq}^{s}(\Omega)$. This completes the proof of i.
Similarly by Proposition 2.13 ii and iii, for every $f\in\widetilde{\mathscr{B}}_{pq}^{s}(\overline{U})$, we have $(f,\phi_{jm}^{g})_{L^{2}}=0$ if $m\notin\mathfrak{S}_{j}$, hence by Proposition 2.13 iii $f=\sum_{j=0}^{\infty}\sum_{m\in\mathfrak{S}_{j},g\in G_{j}}(f,\psi_{jm}^{g})_{L^{2}}\cdot\psi_{jm}^{g}$. Taking $S_{j}\supseteq\mathfrak{S}_{j}\times G_{j}$ be any set in $Q_{j}\times G_{j}$ such that $S_{j}=2^{(j+A)n}$ and applying Proposition 2.13 iv we get $\Lambda:\widetilde{\mathscr{B}}_{pq}^{s}(\overline{U})\to\ell^{q}(2^{j(s-\frac{n}{p}+\frac{n}{2})}\cdot\ell^{p}(S_{j}))_{j=0}^{\infty}$. Applying Lemma 2.15 which gives $E:\mathscr{B}_{pq}^{s}(\Omega)\to\mathscr{B}_{pq}^{s}(\overline{U})$, we get the required boundedness of $\Lambda\circ E$. Since $\Lambda$ and $E$ are both topological embeddings, so does $\Lambda\circ E$. This completes the proof of ii. ∎
@@
Indeed one can check that $\ell^{q_{0}}(2^{j(s_{0}-\frac{n}{p_{0}}+\frac{n}{2})}\cdot\ell^{p_{0}})_{j=0}^{\infty}\subset\ell^{q_{1}}(2^{j(s_{1}-\frac{n}{p_{1}}+\frac{n}{2})}\cdot\ell^{p_{1}})_{j=0}^{\infty}$ holds if and only if either (2.5) holds; and the map is always non-compact. By Proposition 2.13 iv we get i.
Similarly $\ell^{q_{0}}(2^{j(s_{0}-\frac{n}{p_{0}}+\frac{n}{2})}\cdot\ell^{p_{0}}\{1,\dots,2^{nj}\})_{j=0}^{\infty}\subset\ell^{q_{1}}(2^{j(s_{1}-\frac{n}{p_{1}}+\frac{n}{2})}\cdot\ell^{p_{1}}\{1,\dots,2^{nj}\})_{j=0}^{\infty}$ if and only if either (2.6) or (2.7) holds; and the map is compact if and only if (2.7) holds.